Circular No. 5792

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) Astrometric observations have been reported at two observatories on

seven nights since the comet returned to the morning sky in February. Continuation to the ephemeris on IAUC 5715:

1993TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 13 3^{h}47.^{m}8 −70^{o}51.^{′}9 2.267 2.472 89.^{o}4 24.^{o}1 12.2
18 4 28.4 −70 53.8 2.304 2.530 90.8 23.6 12.4
23 5 07.4 −70 33.6 2.348 2.588 91.9 23.0 12.5
28 5 43.63 −69 55.0 2.399 2.646 92.6 22.5 12.7
June 2 6 16.42 −69 02.4 2.456 2.703 92.9 22.0 12.9
7 6 45.65 −68 00.4 2.518 2.760 92.8 21.5 13.1
12 7 11.56 −66 53.0 2.586 2.817 92.4 21.1 13.3
17 7 34.51 −65 43.4 2.659 2.874 91.6 20.7 13.5
22 7 54.92 −64 34.3 2.737 2.930 90.6 20.3 13.7
27 8 13.18 −63 27.5 2.818 2.986 89.4 19.9 13.9
July 2 8 29.62 −62 24.3 2.903 3.042 88.0 19.5 14.1
7 8 44.53 −61 25.6 2.990 3.097 86.4 19.1 14.2
12 8 58.18 −60 31.9 3.080 3.152 84.6 18.7 14.4

CATALOGUE OF COMETARY ORBITS 1993 By popular request, a new edition — the eighth — of the Catalogue

of Cometary Orbits is now available, little more than a year after its predecessor. The 102 pages contain orbital elements (in the J2000.0 system) for 1392 cometary apparitions through April 1993. Greater efficiency in printing now allows the names of the comets to appear together with the orbits. The new catalogue sells at the same price as the seventh edition: $20.00, although subscribers to these Circulars can purchase it for $10.00 and have their accounts debited by this amount. The main part of the catalogue and the table of ‘original’ and ‘future’ 1/a values for the 289 long-period comets with orbits of the highest quality can be supplied by e-mail for $50.00; they are also available on an MS-DOS diskette (5.25-inch or 3.5-inch) for $100.00, this including a facility for extracting individual orbits and computing ephemerides. Checks accompanying orders (or in cases where there is insufficient credit in the accounts for these Circulars) should be made out to “Central Bureau for Astronomical Telegrams” and mailed to the postal address given above.

1993 May 17 Daniel W. E. Green